# Polylogarithmic motivic Chabauty–Kim for $\mathbb{P}^{1}\setminus\{0,1,\infty\}$: the geometric step via resultants

### v2

@@ 6.3. Computation of the Chabauty–Kim function $\mathrm{Res}(\nu_{4}(X),\nu_{6}(X))$
We now compute the polynomials $\nu_{4}(X)$ and $\nu_{6}(X)$ evoked in Section 6.2.3, in order to obtain a more explicit expression for the Chabauty–Kim function $\mathrm{Res}(\nu_{4}(X),\nu_{6}(X))$. Such an expression will be instrumental in proving its non-triviality, and in computing its degree.
@@ 6.3.2. The polynomial $\nu_{4}(X)$
We start with equation (6.16), which is the formula for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{4})$ multiplied by $\Delta$. We inject equation (6.8) to eliminate $\Phi_{e_{1}e_{0}^{2}}^{\sigma_{3}}$. Moreover, whenever $\Delta\Phi_{e_{1}}^{\tau}$ appears, we separate the terms in $\tau=\tau_{p}$ and $\tau=\tau_{q}$. We obtain
@@
Injecting equations (6.18) and (6.22) in equation (6.25) to eliminate $\Phi_{e_{0}}^{\tau_{q}}$, and clearing denominators by multiplying by $f_{\tau_{q}}^{4}f_{\sigma_{3}}$, we obtain an equation of the form $P_{4}(\Phi_{e_{0}}^{\tau_{p}})=0$ where $P_{4}(X)=\theta^{\#}(\nu_{4}(X))$ with
@@ 6.3.3. The polynomial $\nu_{6}(X)$
We start with equation (6.17) which is the formula for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{6})$ multiplied by $\Delta$. We inject (6.8) and (6.11) to eliminate $\Phi_{e_{1}e_{0}^{2}}^{\sigma_{3}}$ and $\Phi_{e_{1}e_{0}^{2}}^{\sigma_{5}}$. We obtain
@@
Injecting equations (6.18) and (6.22) in equation (6.29) to eliminate $\Phi_{e_{0}}^{\tau_{q}}$, and clearing denominators by multiplying by $f_{\tau_{q}}^{6}f_{\sigma_{3}}f_{\sigma_{5}}$, we obtain an equation of the form $P_{6}(\Phi_{e_{0}}^{\tau_{p}})=0$ where $P_{6}(X)=\theta^{\#}(\nu_{6}(X))$ with
@@
[…]
@@ 6.5. Chabauty–Kim functions of higher degree
We have constructed polynomials $P_{4}(X)$ and $P_{6}(X)$, as well as $\nu_{4}(X)$ and $\nu_{6}(X)$, using as starting point the equations (6.7) for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{4})$ and $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{6})$. More generally, a similar process of elimination as in Section 6.2 would give rise, for any $d\geq 2$, to polynomials $P_{2d}(X)\in\theta^{\#}(\mathcal{O}(U_{S})[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{2d}^{\mathfrak{u}}])$ and $\nu_{2d}(X)\in\mathcal{O}(U_{S})[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{2d}^{\mathfrak{u}}]$ such that $P_{2d}(\Phi_{e_{0}}^{\tau_{p}})=0$ and $P_{2d}(X)=\theta^{\#}(\nu_{2d}(X))$. Given $d_{1}>d_{2}\geq 2$, we then obtain polylogarithmic motivic Chabauty–Kim functions
@@
The resultant method for $|S|=2$ thus provides an infinite family of Chabauty–Kim functions. In practice, these other functions will be of high degree and the complexity of the arithmetic step will increase many-fold. [^4] We expect the degree of $\nu_{2d}(X)$ to be $22d-$ 2. In fact, we checked this for $d=2,3,4,5$ in SageMath [sagemath]. We also verified that the leading coefficient of $\nu_{2d}(X)$ is a multiple of $F^{|1|}_{2,2}$ for $d=2,3,4,5$.
@@ 7. Towards the multiple primes case
Let $s=|S|$. We then have $\Sigma_{-1}=\{\tau_{p_{1}},\ldots,\tau_{p_{s}}\}$. The map $\theta^{\#}$ is given by the usual formulas in Definition 4.1. The system of equations is therefore similar to the one in Section 6.1, except that the sums that are indexed by $\Sigma_{-1}$ now have $s$ terms. In particular, the system is still linear with respect to $\Phi^{\sigma_{2n-1}}_{e_{1}e_{0}^{2n-2}}$ and the $\Phi_{e_{1}}^{\tau}$’s, but now the number of variables $\Phi_{e_{0}}^{\tau}$ is $s$, as is the number of variables $\Phi_{e_{1}}^{\tau}$.
A generalisation of the elimination process described in Section 6.2 goes as follows. Take as starting point the equation for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{2d})$ and proceed to eliminate the variables $\Phi^{\sigma_{2n-1}}_{e_{1}e_{0}^{2n-2}}$ with $n\leq d$ by induction on $n$, starting with $n=d$. Then proceed to invert a linear system to eliminate the $\Phi_{e_{1}}^{\tau}$’s, provided that it is invertible. It is unclear whether the determinant of the corresponding matrix will have as nice a shape as the one in Lemma 6.1. After inverting the system, only the $s$ variables $\Phi_{e_{0}}^{\tau}$ remain. In order to eliminate these, proceed by induction on $n$ as follows.

### v3

@@ 6.3. Computation of the Chabauty–Kim function $\mathrm{Res}(\nu_{4}(X),\nu_{6}(X))$
We now compute the polynomials $\nu_{4}(X)$ and $\nu_{6}(X)$ evoked in Section˜6.2.3, in order to obtain a more explicit expression for the Chabauty–Kim function $\mathrm{Res}(\nu_{4}(X),\nu_{6}(X))$. Such an expression will be instrumental in proving its non-triviality, and in computing its degree.
@@ 6.3.2. The polynomial $\nu_{4}(X)$
We start with equation (6.16), which is the formula for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{4})$ multiplied by $\Delta$. We insert equation (6.8) to eliminate $\Phi_{e_{1}e_{0}^{2}}^{\sigma_{3}}$. Moreover, whenever $\Delta\Phi_{e_{1}}^{\tau}$ appears, we separate the terms in $\tau=\tau_{p}$ and $\tau=\tau_{q}$. We obtain
@@
Substituting equations (6.18) and (6.22) in equation (6.25) to eliminate $\Phi_{e_{0}}^{\tau_{q}}$, and clearing denominators by multiplying by $f_{\tau_{q}}^{4}f_{\sigma_{3}}$, we obtain an equation of the form $P_{4}(\Phi_{e_{0}}^{\tau_{p}})=0$ where $P_{4}(X)=\theta^{\#}(\nu_{4}(X))$ with
@@ 6.3.3. The polynomial $\nu_{6}(X)$
We start with equation (6.17) which is the formula for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{6})$ multiplied by $\Delta$. We insert (6.8) and (6.11) to eliminate $\Phi_{e_{1}e_{0}^{2}}^{\sigma_{3}}$ and $\Phi_{e_{1}e_{0}^{2}}^{\sigma_{5}}$. We obtain
@@
Inserting equations (6.18) and (6.22) in equation (6.29) to eliminate $\Phi_{e_{0}}^{\tau_{q}}$, and clearing denominators by multiplying by $f_{\tau_{q}}^{6}f_{\sigma_{3}}f_{\sigma_{5}}$, we obtain an equation of the form $P_{6}(\Phi_{e_{0}}^{\tau_{p}})=0$ where $P_{6}(X)=\theta^{\#}(\nu_{6}(X))$ with
@@
The statement about the degreelwas verified using SageMath [sagemath] (see our GitHub repository for the code). It is then clear from (6.30) that $\nu_{6}(X)$ is a multiple of $\log^{\mathfrak{u}}$. ∎
@@ 6.5. Chabauty–Kim functions of higher degree
We have constructed polynomials $P_{4}(X)$ and $P_{6}(X)$, as well as $\nu_{4}(X)$ and $\nu_{6}(X)$, using as starting point the equations (6.7) for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{4})$ and $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{6})$. More generally, a similar process of elimination as in Section˜6.2 would give rise, for any $d\geq 2$, to polynomials $P_{2d}(X)\in\theta^{\#}(\mathcal{O}(U_{S})[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{2d}^{\mathfrak{u}}])$ and $\nu_{2d}(X)\in\mathcal{O}(U_{S})[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{2d}^{\mathfrak{u}}]$ such that $P_{2d}(\Phi_{e_{0}}^{\tau_{p}})=0$ and $P_{2d}(X)=\theta^{\#}(\nu_{2d}(X))$. Given $d_{1}>d_{2}\geq 2$, we then obtain polylogarithmic motivic Chabauty–Kim functions
@@
The resultant method for $|S|=2$ thus provides an infinite family of Chabauty–Kim functions. In practice, these other functions will be of high degree and the complexity of the arithmetic step will increase many-fold.  [^9] We expect the degree of $\nu_{2d}(X)$ to be 2d22d-2. In fact, we checked this for $d=2,3,4,5$ in SageMath [sagemath]. We also verified that the leading coefficient of $\nu_{2d}(X)$ is a multiple of $F^{|1|}_{2,2}$ for $d=2,3,4,5$.
@@ 7. Towards the multiple primes case
Let $s=|S|$. We then have $\Sigma_{-1}=\{\tau_{p_{1}},\ldots,\tau_{p_{s}}\}$. The map $\theta^{\#}$ is given by the usual formulas in Definition 4.1. The system of equations is therefore similar to the one in Section˜6.1, except that the sums that are indexed by $\Sigma_{-1}$ now have $s$ terms. In particular, the system is still linear with respect to $\Phi^{\sigma_{2n-1}}_{e_{1}e_{0}^{2n-2}}$ and the $\Phi_{e_{1}}^{\tau}$’s, but now the number of variables $\Phi_{e_{0}}^{\tau}$ is $s$, as is the number of variables $\Phi_{e_{1}}^{\tau}$.
A generalisation of the elimination process described in Section˜6.2 goes as follows. Take as starting point the equation for $\theta^{\#}(\operatorname{Li}^{\mathfrak{u}}_{2d})$ and proceed to eliminate the variables $\Phi^{\sigma_{2n-1}}_{e_{1}e_{0}^{2n-2}}$ with $n\leq d$ by induction on $n$, starting with $n=d$. Then proceed to invert a linear system to eliminate the $\Phi_{e_{1}}^{\tau}$’s, provided that it is invertible. It is unclear whether the determinant of the corresponding matrix will have as nice a shape as the one in Lemma 6.1. After inverting the system, only the $s$ variables $\Phi_{e_{0}}^{\tau}$ remain. In order to eliminate these, proceed by induction on $n$ as follows.